OpenStudy (anonymous):

I have the answer can someone just check to see if I'm right. WHich of the following is f '(x) for f(x)=(ln4)^x

OpenStudy (anonymous):

2ln(2)^x(ln(2ln(2))

OpenStudy (anonymous):

Ididn't get that I got (ln)(ln4)^x

OpenStudy (anonymous):

well mine is more simplified

OpenStudy (anonymous):

do u want me to prove the answer?

OpenStudy (anonymous):

it doesn't matter all my choices are multiple choice so if my answer is not correct I guess I don't knwo what the answer is.

OpenStudy (phi):

I got (ln)(ln4)^x <-- this does not quite make sense

OpenStudy (anonymous):

yes i know we use D operator and ur answer would be

OpenStudy (anonymous):

yes the answer is wrong

OpenStudy (anonymous):

ok so then what is the answer I am confussed

OpenStudy (phi):

how about ln(ln(4)) * (ln(4))^x

OpenStudy (phi):

you could write ln(4) as ln(2^2) = 2 ln(2) but only if they put the answer in those terms

OpenStudy (phi):

if you have the problem d/dx a^x write a as e^(ln(a)) and a^x is e^(x ln(a)) in your case a is ln(4)

OpenStudy (anonymous):

so then is the answwer ln(ln(ln4)^x

OpenStudy (anonymous):

no

OpenStudy (phi):

\[ \frac{d}{dx} e^{x ln(ln(4))} \]

OpenStudy (anonymous):

what about this ln(ln4)(ln(ln4))^x

OpenStudy (phi):

the ln(ln(4)) is just a constant (ugly maybe, but a constant) so you get \[ ln(ln(4)) e^{x ln(ln(4))} \] but we can replace \[ e^{x ln(ln(4))} = \left( e^{ln(ln(4))}\right)^x = (ln(4))^x\] and the answer is \[ ln(ln(4))(ln(4))^x \]

OpenStudy (anonymous):

yes that is similar to the one at the top

OpenStudy (anonymous):

so then which is right the one that phi just wrote

OpenStudy (anonymous):

yes phi is right obvious

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

wait thoughthat is not one of my choices

OpenStudy (anonymous):

would ln(ln4)(ln4)^x work